# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

### v1

@@ 3.2 Segmentation
Surface segmentation is a customary preprocessing step in the analysis of individual fragments for reassembly tasks, aimed at breaking down complex surface geometries into simpler, meaningful regions. Typically, this involves distinguishing between the original and fracture surfaces, facilitating easier analysis and subsequent matching with other fragments to achieve successful reassembly.
@@
Geometry-based method. Geometry-based surface segmentation methods rely on the intrinsic geometric properties of the fragment’s surface, such as curvature, edges, and other spatial features. These methods are among the most established techniques in the field due to their reliance on fundamental geometric principles (see Figure 3 (Top)). Several geometry-based methods have been developed for surface segmentation. For instance, [51] employ a simple region-growing algorithm primarily based on normal vectors, along with a facet segmentation threshold, to initially segment the original object into distinct surface regions. Following this segmentation, bumpiness is used to detect fractured surfaces. [52] first perform multi-scale edge extraction on point-sampled surfaces, which is constrained to return closed cycles to obtain initial face cycles. These face cycles are then partitioned into original and fracture surfaces using an iterative normalized cut method on a weighted graph, constructed considering surface sharpness and roughness. [53] use a hierarchical agglomerative clustering algorithm that starts with each element as its own cluster, merging the closest clusters based on a user-defined metric until no further merges are possible. A custom 2-level caching scheme is used to reduce redundant distance calculations and sorting, enhancing efficiency. [54] first apply the Laplace operator to smooth the surface, and then use a clustering algorithm based on a vertex normal vector to acquire rough surface segmentation, remove noise surfaces, and then refine the segmentation by merging adjacent faces using surface roughness and face normal vectors. In addition to these, various other clustering and graph segmentation algorithms, such as K-means clustering [55], hierarchical clustering [56], mean-shift clustering [57], spectral clustering [58] and random walks [59], can also be adapted for fractured surface segmentation. These techniques are well-established, relying on explicit rules and geometric properties. They tend to be more interpretable and less data-intensive but are limited by surface complexity and noise.
@@
Figure 3: (Top) Non-deep learning methods for fracture surface segmentation [51]. (Bottom) Deep learning methods for fracture surface segmentation [60].
@@
Figure 4: A representative multi-piece shape analysis framework presented in [52]. Pair-wise matching use use patch features extracted from fracture surface and segment features extracted from fracture surface boundaries. A voting-based approach is used to find consistent correspondences between geometric features. Matching among pieces employ a greedy approach.
@@
Learning-based method. Learning-based surface segmentation techniques have gained traction in recent years due to their ability to automatically learn patterns and features from data, without the need for explicit geometric rules. These methods, particularly those involving deep learning, are transforming how surface segmentation is performed in fragment reassembly tasks. Although some learning-based fractured object reassembly methods take the entire fragments as input for analysis, bypassing prior surface segmentation, there are still methods that decouple the reassembly pipeline, retaining segmentation to achieve better matching and reduce computational costs. Learning-based segmentation methods are predominantly based on graph learning techniques. The initial segmentation process can be conceptualized as a binary classification task. For example, [60] utilize a convolutional neural network (CNN) to learn the labels of ridge and non-ridge cells (see Figure 3(Bottom)), facilitating the extraction of the fracture margins from the fragments. Subsequently, fractured surface can be well segmented by trimming the fragment models along these margins. [61] employ Multi-Layer Perceptron (MLP) layers combined with a sigmoid activation function to predict the confidence in identifying a fracture point. If we generalize segmenting fractured surfaces to 3D semantic segmentation, we can also leverage advanced general-purpose graph learning algorithms like PointNet [62], PointNet++ [63] and DGCNN [64] to perform segmentation. These powerful frameworks offer robust solutions for intricate 3D data analysis, and they will also be mentioned in the next section. These approaches, particularly deep learning, offer more flexibility and can handle complex surfaces and noisy data better. However, they require large datasets and significant computational power.
@@ 4.1 Pairwise Matching
Rigid scan maching. A relevant problem of pairwise piece matching is rigid scan matching, which finds the relative rigid pose between two depth scans. Rigid scan matching is an important step in 3D geometry reconstruction from depth scans. A standard pipeline of rigid scan matching consists of feature extraction, feature matching, and rigid pose fitting. Early approaches use hand-crafted geometric descriptors, e.g., SpinImage [65] and Integral invariant [66] features. These features usually only encode local information; therefore, the resulting feature matches have a lot of false positives. The rigid pose fitting procedure addresses this issue by enforcing the rigidity constraint, which finds a subset of feature matches that share a rigid transformation. Popular rigid pose fitting approaches include RANSAC [67], general Hough transform [68, 69, 70], and their variants [71, 72]. Another category of methods formulates rigid scan matching as spectral matching, which computes a consistency matrix among candidate feature matches. The elements of this consistency matrix encode whether distances and angles are preserved between the corresponding pairs of correspondences. Rigid pose fitting is formulated as computing a sub-matrix in which all elements are consistent, and this is done using spectral techniques.
More recent approaches develop deep learning techniques for rigid scan matching. These approaches employ deep neural networks to extract features and develop differentiable layers for fitting rigid poses. Examples include differentiable RANSAC and attention-based rigid regression. Representative methods in this category include DCP [73], PNetLK [74], FCGF [75], Predator [76], RoReg [77], and GeoTF [78]. These approaches show significantly improved performance on benchmark datasets where the domain gap between the testing set and the training set is small. However, it is still an open question to develop generalizable deep learning approaches for rigid scan matching.
Non deep fragment matching. Similar approaches have been developed for rigid fragment matching, which also combines feature extraction, feature matching, and rigid pose fitting. However, rigid fragment matching and rigid scan matching have several notable differences. These differences come from several aspects, including where to extract features and what constraints are induced from feature matching. Another important difference is the robustness of the extracted features, the feature descriptors, and the matching procedure, as the surfaces of the fragments may be severely damaged or there are tiny fragments between two fragments, causing geometric differences. Figure 4 shows a representative framework for fragment matching.
@@
Most approaches use features extracted from the fracture surfaces. This requires fracture surface segmentation as discussed in Section 3.2. Robustness has been a major concern when extracting features from fracture surfaces. Therefore, many features introduced in the literature are patch features that are introduced to improve robustness. [52] use patches defined by sub-level sets of the integral invariant descriptors and summarize geometric descriptors derived from performing principal component analysis of patch shapes. [81] introduced an approach to smooth the fragments and segment the concave and convex regions of the fracture surfaces for robust matching. [82] studied multi-scale covariance matrix based point descriptors. [83] introduced an algorithm based on cluster-tree tailored for matching fragments with salient fractured surfaces. A key issue in matching fracture surfaces is developing descriptors that can be used to identify geometric patches that can be tucked together [84]. A recent paper [61] also addresses this issue by introducing a primal-dual descriptor in deep learning approach.
@@
Many papers focus on extract from boundary curves of fracture surfaces [84, 81]. Such features are more robust in the setting where the areas of the fracture surfaces are small, making it difficult to compute and extract geometric descriptors. [85] analyze and simulate fracture patterns of Theran wall paintings. It builds statistical models of boundary curves of wall painting fragments, which can be applied to match boundary curves. [84] detect boundary curves based on curvature computation and apply FFT to compute the descriptors of points on the boundary curves for fragment matching. [86] introduced a multi-channel boundary contour representation for matching sub-contours via 2D partial image registration.[87] perform curvature computation to detect parabolic contours, whose descriptors are determined by moments. [88] studied how to detect boundary curves of thin fragments. Fragment matching is done using voting. [89] introduced a fast algorithm for 2D fragment assembly based on a partial EMD distance. [90] introduced a descriptor based on the convex/concave information of a point on the boudnary curves of the fractured surface. The similarity between two fractured surfaces is calculated based on the spin images of the point of the feature curve and the distance and normal deviation between the two feature curves to find the matching fractured surface. [91] utilized the detected breaking curves as cues for fracture surface segmentation and registration. [92] introduced a new matching algorithm for stone tool reassembly based on detecting and matching contour points on flake surfaces. Matching is done using a five-point method, which extends Super4PCS. For efficiency concerns, it focuses on matching flake surfaces to a core stone, which is applicable in this specific setting. In [93], a set of matching units is detected and described by the 2D Link-Chain Descriptors (LCD) and the 3D Spatial-Distribution Descriptors (SDD). Second, the pairwise reassembly probability is calculated using LCD and SDD descriptors; then, collision detection is performed to eliminate the incorrect overlapping pairs.
@@
Several approaches rely on features on the original surfaces for matching (see Figure 5). Early work by [94] discussed how to use original surface features for geometric matching. Examples include the continuity of texture patterns or matches between curve features extracted from original surfaces. [86] introduced a similar approach to enforce the original surface consistency. The former approach can be found in most reassembly pipelines of wall painting fragments, e.g., [79]. [80] studied how to match fragments that are under erosion and proposed an approach that completes the original surfaces of the fragments and matches the completed surfaces. In this approach, the original non-overlapping surfaces become overlapping, allowing us to detect and match extracted feature points.
@@
In many cases, the approach of extracting and matching features is not robust when fragments are undergoing serious erosion, which is the case for most real datasets. In [96], the authors find that detecting and matching feature points from fractured surfaces is much more challenging than matching scans. They proposed confidence scores on potential matches. Several approaches define scoring functions over the space of transformations and develop global search strategies to global optimal solutions. [97] introduced an approach that focuses on wall-painting fragments and applies a discrete search of relative poses to identify global solutions to reassembly. [95] introduced four criteria for fragment matching (see Figure 6). The first criterion exploits the volume of the gap between two properly placed fragments. The second considers the overlap of the fragments in each possible matching position. Criteria 3 and 4 employ principles from the calculus of variations to obtain bounds for the area and the mean curvature of the contact surfaces and the length of contact curves, which must hold if the two fragments match.
In addition to geometric matching, another difference between scan matching and fragment matching is that the matches should be free of penetrations. This constraint is easy to integrate for RANSAC-based approaches [67, 98]. An open question is how to enforce this under global optimization approaches such as spectral matching. Regarding local refinement methods, several papers [52, 99, 100, 101] studied how to formulate constrained optimization techniques to enforce non-penetration constraints (See Figure 7 (Left)). [50] introduced a gradient field to optimize the packing of a collection of objects, enforcing the non-penetration constraint (See Figure 7 (Right)).
@@
Deep fragment matching. In the deep learning era, the research community has studied developing deep neural networks to solve the multi-piece reassembly problem. Early work includes Deepzzle [103], which studied image reassembly, a special case of fractured object reassembly. Deepzzle studied how to handle noisy inputs, e.g., erosion, missing pieces, and fragments of other objects. Thanks to the emergence of 3D benchmark datasets for fractured object reassembly, there are increasing interests in developing deep learning approaches to solve the 3D fractured object reassembly. [104] proposed MatchMakerNet (see Figure 9(Left)), a network architecture designed to automate the pairing of object fragments for reassembly. The approach takes two point clouds as input and its network leverages graph convolution alongside a simplified version of DGCNN for prediction.
However, most existing approaches have focused on feature learning and incorporating rigidity constraints and problem-specific knowledge deserves future research.
@@ 4.2 Multi-Piece Matching
Optimization-based based approaches. The difference between multiple-piece matching and pairwise matching is that there is a cycle-consistency constraint [105, 106, 107] among cycles of fragments. Specifically, when composing relative poses along a cycle of fragments, we should arrive at the identity map. If not, then at least one of the relative poses is incorrect. This cycle-consistency constraint can be used to prune incorrect pair-wise matches computed between pairs of objects in isolation.
@@
Another category of approaches formulate multi-piece matching as Markov-random field (MRF) inference. This is formulated as sampling a discrete set of candidate poses of each object and model pairwise matches as pairwise potentials to optimize the best candidate of each object. This approach was first introduced to solve the Jigsaw puzzle problem in images [110, 111], where the candidate positions of each image patch are given by the grid of image patches of a given image. A relevant approach is [112], which presents a linear programming (LP) relaxation to the MRF formulation on the image-based square patch puzzle problem. This MRF formulation was later applied for in MRF-SFM [113, 114] to solve the structure-from-motion problem. The nice properties of this MRF formulation is that it possesses better solvers that offer improved solutions compared to greedy approaches. Moreover, it is easy to incorporate the non-penetration constraint, i.e., if the a pair of fragments penetration under two candidate poses, then the corresponding pairwise potential is infinite. The limitation of this MRF formulation is that it requires many pose samples to obtain accurate solutions.
Modern multi-piece matching typically formulates low-rank matrix recovery by establishing the equivalence between the cycle-consistency constraint and the fact that the matrix that encodes pairwise maps in block is low-rank, c.f. [115]. The optimization approaches fall into categories of semidefinite programming relaxation [116, 117, 115, 118, 119, 120, 121], spectral techniques [122, 123, 107, 124, 125, 126, 127, 128, 129, 130, 131, 132], and non-convex optimization [133, 134, 135, 136, 137, 138, 139]. Their advantages are superior empirical performance and theoretical guarantees. However, it is not easy to enforce the penetration-free constraints under them.
@@
Applications of these multi-object matching approaches in the context multi-fragment reassembly have been limited. Most approaches follow greedy approaches [52, 143, 141]. There are a few exceptions. [144] introduced an unsupervised genetic algorithm to reassemble wall painting fragments. The approach evolves a pool of partial reconstructions that grow through recombination and selection over the course of generations. The authors introduced a novel algorithm for combining partial reconstructions that is robust to noise and outliers, and a new selection procedure that balances fitness and diversity in the population. [86] detect loop in the pose graph and performs loop closure to rectify pose errors. [140] enforce global Consistency in the Automatic assembly of fragmented artifacts. This approach addresses error accumulation problem the global reassembly of many fragments, among which there are many potential matches with a pair of fragments. [93] introduced a global optimal reassembly solution is obtained by iterative graph optimization with the constrains of the loop-closures and overlap restrictions. One critical difference between multi-scan matching and multi-fragment matching is that the underlying graph of the fragments is much more sparse than the graph in multi-scan matching, as scans have significantly overlap, while fragments do not overlap.
@@
Deep pose regression based approaches. The emergence of large-scale datasets such as Breaking Bad [145] has simulated a wave in developing deep learning approaches that perform pose regression. Many approaches adopt recent advances in 3D feature learning using Transformers. [146] developed Transfomers to find pairs of matching fragments and their relations. Specifically, they first employ a graph-based network to extract the local contour and texture features of the fragments. Then, for the pair-searching task, they adopt a linear transformer-based module to integrate these local features and use contrastive loss to encode the global features of each fragment. For the pair-matching task, they design a weighted fusion module to dynamically fuse extracted local contour and texture features, and formulate a similarity matrix for each pair of fragments to calculate the matching score and infer the adjacent segment of contours. [147] introduced Proxy Match Transform (PMT), an approximate high-order feature transform layer that enables reliable correspondences between the mating surfaces of parts while being efficient. Using PMT. they introduced a new framework, dubbed Proxy Match TransformeR (PMTR), for the geometric assembly task. [148] introduced a Transformer-based encoder to extract both global and local geometric features, which are then passed to two decoders to estimate the relative pose between two fragments. [141] introduced a recurrent approach that aligns and merges fragments into larger groups. The key contribution is a diffusion model simultaneously denoises the 6-DoF alignment parameters of the fragments, and a transformer model verifies and merges pairwise alignments into larger ones, whose process is repeated iteratively. [149] introduced an approach that identifies critical stones among the input fragments for efficient computation of the reassembly procedure. This is implemented using the attention mechanism in an iterative manner. [150] use a hybrid attention module to model and reason complex structural relationships between fragment patches. The model has intra- and inter-attention layers, enabling capturing of crucial contextual information within fragments and relative structural knowledge across fragments. The model is enhanced with an adjacency aware hierarchical pose estimator.
@@
People have developed equivalent networks [151] for point cloud learning to solve fractured object reassembly. [142] developed equivariant and invariant encoders of each part to compute equivariant and invariant correlation modules to aggregate extracted signals across all parts, which are used to regress the pose of each input part. [152] introduced an approach, called SE(3)-bi-equivariant transformer (BITR), based on the SE(3)-bi-equivariance prior to the task. Specifically, BITR first extracts features of the inputs using a novel SE(3)×SE(3)-transformer, and then projects the learned feature to group SE(3) as output. BITR can not only handle non-overlapped PCs, but also guarantee robustness against initial positions. Extension of SE(3) equivariance. [153] introduced DiffAssemble, which follows Diffusion Probabilistic models formulations. DiffAssemble models a Markov chain in which it injects noise into the position and orientation of the pieces. The correct and random poses at the initial time step and the final time step, respectively. The denoising network is an attention-based GNN that takes as input a graph where each node contains an equivariant feature that describes a particular piece and its position and orientation. Similarly, [154] and [141] treated the reassembly task as a denoising task and introduced FragmentDiff and PuzzleFusion++, a method that applies diffusion denoising through transformer to predict the pose parameters of each fragment based on their global feature correlations and learned pose priors.

### v2

@@ 3.2 Segmentation
Surface segmentation is a common preprocessing step in the analysis of individual fragments for reassembly tasks, aimed at breaking down complex surface geometries into simpler, meaningful regions. Typically, this involves distinguishing between the original and fracture surfaces, facilitating easier analysis and subsequent matching with other fragments to achieve successful reassembly.
@@
Geometry-based methods. Geometry-based surface segmentation methods rely on the intrinsic geometric properties of the fragment’s surface, such as curvature, edges, and other spatial features. These methods are among the most established techniques in the field due to their reliance on fundamental geometric principles (see Figure 4 (Top)). Several geometry-based methods have been developed for surface segmentation. For example, [52] employs a simple region-growing algorithm based primarily on normal vectors, along with a facet segmentation threshold, to initially segment the original object into distinct surface regions. Following this segmentation, bumpiness is used to detect fractured surfaces. [53] first perform multi-scale edge extraction on point-sampled surfaces, which is constrained to return closed cycles to obtain initial face cycles. These face cycles are then partitioned into original and fracture surfaces using an iterative normalized cut method on a weighted graph, constructed considering surface sharpness and roughness. [54] use a hierarchical agglomerative clustering algorithm that starts with each element as its own cluster, merging the closest clusters based on a user-defined metric until no further merges are possible. A custom 2-level caching scheme is used to reduce redundant distance calculations and sorting, enhancing efficiency. [55] first apply the Laplace operator to smooth the surface, and then use a clustering algorithm based on a vertex normal vector to acquire rough surface segmentation, remove noise surfaces, and then refine the segmentation by merging adjacent faces using surface roughness and face normal vectors. In addition to these, various other clustering and graph segmentation algorithms, such as K-means clustering [56], hierarchical clustering [57], mean-shift clustering [58], spectral clustering [59] and random walks [60], can also be adapted for fractured surface segmentation. These techniques leverage explicit geometric rules and properties, making them interpretable and efficient with limited data. However, their performance may decrease when handling complex surfaces or noisy input.
@@
Figure 4: (Top) Procedural methods for fracture surface segmentation [52]. (Bottom) Deep learning methods for fracture surface segmentation [61].
@@
Figure 5: A representative multi-piece shape analysis framework presented in [53]. Pair-wise matching use patch features extracted from the fracture surface and segment features extracted from fracture surface boundaries. A voting-based approach is used to find consistent correspondences between geometric features. Matching among pieces employs a greedy approach.
@@
Learning-based methods. Learning-based surface segmentation techniques have emerged as a promising direction, as they can automatically discover patterns and features from data without requiring hand-crafted geometric rules. These methods offer new possibilities for surface segmentation in fragment reassembly tasks. Although some learning-based fractured object reassembly methods take the entire fragments as input for analysis, bypassing prior surface segmentation, there are still methods that decouple the reassembly pipeline, retaining segmentation to achieve better matching and reduce computational costs. Learning-based segmentation methods are predominantly based on graph learning techniques. The initial segmentation process can be conceptualized as a binary classification task. For example, [61] utilize a convolutional neural network (CNN) to learn the labels of the ridge and non-ridge cells (see Figure 4(Bottom)), facilitating the extraction of the fracture margins of the fragments. Subsequently, fractured surfaces can be well segmented by trimming fragment models along these margins. [62] employ Multi-Layer Perceptron (MLP) layers combined with a sigmoid activation function to predict confidence in the identification of a fracture point. If we generalize segmenting fractured surfaces to 3D semantic segmentation, we can also leverage advanced general-purpose graph learning algorithms like PointNet [63], PointNet++ [64] and DGCNN [65] to perform segmentation. These powerful frameworks offer robust solutions for intricate 3D data analysis, and they will also be mentioned in the next section. These frameworks provide effective tools for analyzing complex 3D data, as we will discuss further in the next section. While these approaches can potentially handle challenging cases with complex surfaces and noise, they typically require large training datasets and substantial computational resources.
@@ 4.1 Pairwise Matching
Rigid scan matching. A relevant problem of pairwise piece matching is rigid scan matching, which finds the relative rigid pose between two depth scans. Given two point sets $P$ and $Q$, rigid scan matching aims to find a rotation $R\in SO(3)$ and translation $t\in\mathbb{R}^{3}$ that minimize:
$\min_{R,t}\sum_{i=1}^{n}\min_{j}||Rp_{i}+t-q_{j}||^{2}$ (1)
where $p_{i}\in P$ and $q_{j}\in Q$ are points from the respective scans. A standard pipeline of rigid scan matching consists of feature extraction, feature matching, and rigid pose fitting. Early approaches use hand-crafted geometric descriptors, e.g., SpinImage [66] and Integral invariant [67] features. These features usually only encode local information; therefore, the resulting feature matches have many false positives. The rigid pose fitting procedure addresses this issue by enforcing the rigidity constraint, which finds a subset of feature matches that share a rigid transformation. Popular rigid pose fitting approaches include RANSAC [68], general Hough transform [69, 70, 71], and their variants [72, 73]. Another category of methods formulates rigid scan matching as spectral matching [74], which computes a consistency matrix among candidate feature matches. The elements of this consistency matrix encode whether distances and angles are preserved between the corresponding pairs of correspondences. Rigid pose fitting is formulated as computing a sub-matrix in which all elements are consistent, and is done using spectral techniques.
More recent approaches develop deep learning techniques for rigid scan matching. These approaches employ deep neural networks to extract features and develop differentiable layers to fit rigid poses. Examples include differentiable RANSAC and attention-based rigid regression. Representative methods in this category include DCP [75], PNetLK [76], FCGF [77], Predator [78], RoReg [79], and GeoTF [80]. These approaches demonstrate promising results on benchmark datasets with similar training and testing conditions. Developing approaches that generalize well to diverse real-world scenarios remains an important research direction.
Procedural fragment matching. Similar approaches have been developed for rigid fragment matching, which also combines feature extraction, feature matching, and rigid pose fitting. However, rigid fragment matching and rigid scan matching have several notable differences. These differences come from several aspects, including where to extract features and what constraints are induced by feature matching. Another important difference is the robustness of the extracted features, the feature descriptors, and the matching procedure, as the surfaces of the fragments may be severely damaged or there are tiny fragments between two fragments, causing geometric differences. Figure 5 shows a representative framework for fragment matching.
@@
Most approaches use features extracted from the fracture surfaces. This requires fracture surface segmentation as discussed in Section 3.2. Robustness has been a major concern when extracting features from fracture surfaces. Therefore, many features introduced in the literature are patch features that are developped to improve robustness. [53] use patches defined by sub-level sets of the integral invariant descriptors and summarize geometric descriptors derived from performing principal component analysis of patch shapes. [83] introduced an approach to smooth the fragments and segment the concave and convex regions of the fracture surfaces for robust matching. [84] studied multi-scale covariance matrix based point descriptors. [85] introduced an algorithm based on cluster-tree tailored for matching fragments with salient fractured surfaces. A key issue in matching fracture surfaces is developing descriptors that can be used to identify geometric patches that can be tucked together [86]. A recent paper [62] also addresses this issue by introducing a primal-dual descriptor in the deep learning approach.
@@
Many papers focus on extracting boundary curves of fracture surfaces [86, 83]. Such features are more robust in the setting where the areas of the fracture surfaces are small, making it difficult to compute and extract geometric descriptors. [87] performs (potentially partial) contour matching of the break boundaries to constrain the pairwise surface registration. [88] analyze and simulate fracture patterns of Theran wall paintings. It builds statistical models of boundary curves of wall painting fragments, which can be applied to match boundary curves. [86] detect boundary curves based on curvature computation and apply FFT to compute the descriptors of points on the boundary curves for fragment matching. [89] introduced a multi-channel boundary contour representation for matching sub-contours via 2D partial image registration.[90] perform curvature computation to detect parabolic contours, whose descriptors are determined by moments. [91] studied how to detect boundary curves of thin fragments. Fragment matching is done using voting. [92] introduced a fast algorithm for 2D fragment assembly based on a partial EMD distance. [93] introduced a descriptor based on the convex/concave information of a point on the boundary curves of the fractured surface. The similarity between two fractured surfaces is calculated based on the spin images of the point of the feature curve and the distance and normal deviation between the two feature curves to find the matching fractured surface. [94] utilized the detected breaking curves as cues for fracture surface segmentation and registration. [95] introduced a new matching algorithm for stone tool reassembly based on detecting and matching contour points on flake surfaces. Matching is done using a five-point method, which extends Super4PCS. For efficiency concerns, it focuses on matching flake surfaces to a core stone, which is applicable in this specific setting. In [96], a set of matching units is detected and described by the 2D Link-Chain Descriptors (LCD) and the 3D Spatial-Distribution Descriptors (SDD). Second, the pairwise reassembly probability is calculated using LCD and SDD descriptors; then, collision detection is performed to eliminate the incorrect overlapping pairs.
@@
Several approaches rely on the features on the original surfaces for matching (see Figure 6). Early work by [97] discussed how to use original surface features for geometric matching. Examples include the continuity of texture patterns or matches between curve features extracted from original surfaces. [89] introduced a similar approach to enforce the original surface consistency. The former approach can be found in most reassembly pipelines of wall painting fragments, e.g., [81]. [82] studied how to match fragments that are under erosion and proposed an approach that completes the original surfaces of the fragments and matches the completed surfaces. In this approach, the original non-overlapping surfaces become overlapping, allowing us to detect and match extracted feature points.
@@
In many cases, the approach of extracting and matching features is not robust when fragments are undergoing serious erosion, which is the case for most real datasets. In [99], the authors find that detecting and matching feature points from fractured surfaces is much more challenging than matching scans. They proposed confidence scores on potential matches. Several approaches define scoring functions over the space of transformations and develop global search strategies for global optimal solutions. [100] uses simulated annealing to search the space of the source to transformation, while employing hardware-accelerated rasterization to extract surface distances based on depth maps and compute a similarity score using distance gradients. [101] introduced an approach that focuses on wall-painting fragments and applies a discrete search of relative poses to identify global solutions to reassembly. [98] introduced four criteria for fragment matching (see Figure 7). The first criterion exploits the volume of the gap between two properly placed fragments. The second considers the overlap of the fragments in each possible matching position. Criteria 3 and 4 employ principles from the calculus of variations to obtain bounds for the area and the mean curvature of the contact surfaces and the length of contact curves, which must hold if the two fragments match.
In addition to geometric matching, another difference between scan matching and fragment matching is that the matches should be free of penetrations. This constraint is easy to integrate for RANSAC and stochastic global optimization approaches [68, 102]. An open question is how to enforce this under global optimization approaches such as spectral matching. Regarding local refinement methods, several papers [53, 103, 104, 105] studied how to formulate constrained optimization techniques to enforce non-penetration constraints (See Figure 8 (Left)). [51] introduced a gradient field to optimize the packing of a collection of objects, enforcing the non-penetration constraint (See Figure 8 (Right)).
@@
Deep fragment matching. In the deep learning era, the research community has studied developing deep neural networks to solve the multi-piece reassembly problem. Early work includes Deepzzle [107], which studied image reassembly, a special case of fractured object reassembly. Deepzzle explored methods to handle real-world challenges such as erosion, missing pieces, and the presence of unrelated fragments. Due to the emergence of 3D benchmark datasets for fractured object reassembly, there is increasing interest in developing deep learning approaches to solve the 3D fractured object reassembly. [108] proposed MatchMakerNet (see Figure 10(Left)), a network architecture designed to automate the pairing of object fragments for reassembling. The approach takes two point clouds as input and its network leverages graph convolution alongside a simplified version of DGCNN for prediction. Current approaches have primarily focused on feature learning while incorporating rigidity constraints and domain-specific knowledge remains an important direction for future research. Bridging deep learning methods with practical applications continues to be an active area of investigation.
@@ 4.2 Multi-Piece Matching
Optimization-based based approaches. The difference between multiple-piece matching and pairwise matching is the cycle-consistency constraint [109, 110, 111] among cycles of fragments. Specifically, when composing relative transformations $T_{ij}$ along a cycle of fragments $(i_{1},i_{2},...,i_{k})$, we should have:
$T_{i_{1}i_{2}}T_{i_{2}i_{3}}...T_{i_{k-1}i_{k}}T_{i_{k}i_{1}}=I$ (2)
where $I$ is the identity transformation. If this constraint is violated, then at least one of the relative poses along the cycle must be incorrect. This cycle-consistency constraint is used to prune incorrect pair-wise matches computed between pairs of objects in isolation.
@@
Another category of approaches formulates multi-piece matching as Markov-random field (MRF) inference. This is formulated as sampling a discrete set of candidate poses of each object and model pairwise matches as pairwise potentials to optimize the best candidate of each object. This approach was first introduced to solve the Jigsaw puzzle problem in images [114, 115], where the candidate positions of each image patch are given by the grid of image patches of a given image. A relevant approach is [116], which presents a linear programming (LP) relaxation to the MRF formulation on the image-based square patch puzzle problem. This MRF formulation was later applied in MRF-SFM [117, 118] to solve the structure-from-motion problem. The nice properties of this MRF formulation are that it possesses better solvers that offer better solutions compared to greedy approaches. Moreover, it is easy to incorporate the non-penetration constraint, i.e., if two fragment penetrates under two candidate poses, then the corresponding pairwise potential is infinite. The limitation of this MRF formulation is that it requires many pose samples to obtain accurate solutions.
Modern multi-piece matching typically formulates low-rank matrix recovery by establishing the equivalence between the cycle-consistency constraint and the fact that the matrix that encodes pairwise maps in the block is low-rank, c.f., [119]. The optimization approaches fall into categories of semidefinite programming relaxation [120, 121, 119, 122, 123, 124, 125], spectral techniques [126, 127, 111, 128, 129, 130, 131, 132, 133, 134, 135, 136], and non-convex optimization [137, 138, 139, 140, 141, 142, 143]. Their advantages are superior empirical performance and theoretical guarantees. However, it is not easy to enforce the penetration-free constraints under them.
@@
Applications of these multi-object matching approaches in the context of multi-fragment reassembly have been limited. Most approaches follow greedy approaches [53, 147, 145]. There are a few exceptions. [148] introduced an unsupervised genetic algorithm to reassemble fragments of wall paintings. The approach evolves a pool of partial reconstructions that grow through recombination and selection over the course of generations. The authors introduced a novel algorithm for combining partial reconstructions that is robust to noise and outliers, and a new selection procedure that balances fitness and diversity in the population. [89] detect loop in the pose graph and perform loop closure to rectify pose errors. [144] enforce global Consistency in the Automatic assembly of fragmented artifacts. This approach addresses error accumulation problem the global reassembly of many fragments, among which there are many potential matches with a pair of fragments. [96] introduced a global optimal reassembly solution is obtained by iterative graph optimization with the constrains of the loop-closures and overlap restrictions. One critical difference between multi-scan matching and multi-fragment matching is that the underlying graph of the fragments is much more sparse than the graph in multi-scan matching, as scans have significant overlap, while fragments do not overlap.
@@
Deep pose regression based approaches. The emergence of large-scale datasets such as Breaking Bad [149] has simulated a wave in developing deep learning approaches that perform pose regression. Many approaches adopt recent advances in 3D feature learning using Transformers. [150] developed Transfomers to find pairs of matching fragments and their relations. Specifically, they first employ a graph-based network to extract the local contour and texture features of the fragments. Then, for the pair-searching task, they adopt a linear transformer-based module to integrate these local features and use contrastive loss to encode the global features of each fragment. For the pair-matching task, they design a weighted fusion module to dynamically fuse extracted local contour and texture features, and formulate a similarity matrix for each pair of fragments to calculate the matching score and infer the adjacent segment of contours. [151] introduced Proxy Match Transform (PMT), an approximate high-order feature transform layer that enables reliable correspondences between the mating surfaces of parts while being efficient. Using PMT. they introduced a new framework, dubbed Proxy Match TransformeR (PMTR), for the geometric assembly task. [152] introduced a Transformer-based encoder to extract both global and local geometric features, which are then passed to two decoders to estimate the relative pose between two fragments. [145] introduced a recurrent approach that aligns and merges fragments into larger groups. The key contribution is a diffusion model that simultaneously denoises the 6-DoF alignment parameters of the fragments, and a transformer model that verifies and merges pairwise alignments into larger ones, whose process is repeated iteratively. [153] introduced an approach that identifies critical stones among the input fragments for efficient computation of the reassembly procedure. This is implemented using the attention mechanism in an iterative manner. [154] use a hybrid attention module to model and reason complex structural relationships between fragment patches. The model has intra- and inter-attention layers, enabling the capturing of crucial contextual information within fragments and relative structural knowledge across fragments. The model is enhanced with an adjacency aware hierarchical pose estimator.
@@
Researchers have developed equivalent networks [155] for point cloud learning to solve fractured object reassembly. [146] developed equivariant and invariant encoders of each part to compute equivariant and invariant correlation modules to aggregate extracted signals across all parts, which are used to regress the pose of each input part. [156] introduced an approach, called SE(3)-bi-equivariant transformer (BITR), based on the SE(3)-bi-equivariance prior to the task. Specifically, BITR first extracts features of the inputs using a novel SE(3)×SE(3)-transformer, and then projects the learned feature to group SE(3) as output. BITR can not only handle non-overlapped PCs, but also guarantee robustness against initial positions. [157] introduced DiffAssemble, which follows Diffusion Probabilistic model formulations. DiffAssemble models a Markov chain in which it injects noise into the position and orientation of the pieces. The correct and random poses at the initial time step and the final time step, respectively. The denoising network is an attention-based GNN that takes as input a graph where each node contains an equivariant feature that describes a particular piece and its position and orientation. Similarly, [158] and [145] treated the reassembly task as a denoising task and introduced FragmentDiff and PuzzleFusion++, methods that apply diffusion denoising through the Transformers to predict the pose parameters of each fragment based on their global feature correlations and learned pose priors. While these deep learning approaches show promising results, bridging the gap between synthetic training data and real-world applications remains an important challenge. Real fragments often exhibit erosion, weathering, or missing pieces - conditions that are difficult to fully capture in synthetic training data. Understanding how to better generalize to these practical scenarios is an active area of research.